# EFSA: Towards Event-Level Financial Sentiment Analysis

### v1

@@ Abstract
In this paper, we extend financial sentiment analysis (FSA) to event-level since events usually serve as the subject of the sentiment in financial text. Though extracting events from the financial text may be conducive to accurate sentiment predictions, it has specialized challenges due to the lengthy and discontinuity of events in a financial text. To this end, we reconceptualize the event extraction as a classification task by designing a categorization comprising coarse-grained and fine-grained event categories. Under this setting, we formulate the Event-Level Financial Sentiment Analysis (EFSA for short) task that outputs quintuples consisting of (company, industry, coarse-grained event, fine-grained event, sentiment) from financial text. A large-scale Chinese dataset containing $12,160$ news articles and $13,725$ quintuples is publicized as a brand new testbed for our task. A four-hop Chain-of-Thought LLM-based approach is devised for this task. Systematically investigations are conducted on our dataset, and the empirical results demonstrate the benchmarking scores of existing methods and our proposed method can reach the current state-of-the-art. Our dataset and framework implementation are available at https://anonymous.4open.science/r/EFSA-645E
@@
Tianyu Chen1,2111These authors contributed equally to this work., Yiming Zhang3111These authors contributed equally to this work., Guoxin Yu1,2, Dapeng Zhang4,Li Zeng6, Qing He1,2,3, Xiang Ao1,2,5222Corresponding author. 1Key Laboratory of AI Safety and Security, Chinese Academy of Sciences(CAS), Institute of Computing Technology, CAS, Beijing 100190, China. 2University of Chinese Academy of Sciencs, Beijing 100049, China 3Henan Institute of Advanced Technology, Zhengzhou University, Zhengzhou 450002, P.R. China 4School of IoT Engineering, Jiangsu Vocational College of Information Technology, Wuxi, 214153, China; 5Institute of Intelligent Computing Technology, Suzhou, CAS 6Information Technology Department I, Shenzhen Stock Exchange {chentianyu22s,aoxiang}@ict.ac.cn
@@ 1 Introduction
To this end, we aim to discover the events in financial text to provide an easier financial sentiment prediction. An intuitive way to identify an event from financial text is extracting a specific text span from the original text just like existing aspect-based sentiment analysis (ABSA) that finds both aspect and opinion terms from customer’s comments (Pontiki et al., 2016; Zhang et al., 2022; Yu et al., 2023). However, direct adapting approaches of ABSA for this task could be ineffective even infeasible due to the events in the financial text being overlong and discontinuous (c.f. Figure 1). We will provide a further discussion about the similarities and differences between ABSA and our task in Section 2.3.
In this paper, based on our observations, we provide an alternative setting for FSA and propose a novel task, named Event-Level Financial Sentiment Analysis (EFSA), involving the prediction of quintuples (company, sector, coarse-grained event, fine-grained event, sentiment). Here we have enhanced the FSA tasks in two facets. First, to overcome the difficulties associated with extracting events from financial texts, we reconceptualize the event extraction task as a classification task. We design a categorization comprising coarse-grained and fine-grained event categories specifically for various events in financial news. Besides, we construct a knowledge-based rule to classify companies by industry, enabling FSA to elevate to higher dimensions, such as indices or sector markets. Our task setting can offer substantial practical value in financial applications, including stock trading, stock market anomaly attribution analysis, enterprise risk management, etc. Figure 1 presents example quintuples in our EFSA task.
To support this task, we annotated a large-scale dataset from Chinese financial news. The dataset includes $12,160$ news articles, selected from an initial set of over $50,000$ articles collected from mainstream Chinese financial news websites. Detailed annotations were conducted based on the above task settings. To the best of our knowledge, this dataset is a large-scale fine-grained annotated dataset for FSA and the largest Chinese dataset in the event-level financial sentiment analysis domain.
@@
  • We conduct systemically benchmark experiments to investigate the efficacy of existing methods for the EFSA task and introduce a novel LLM-based framework reaching the current state-of-the-art.
@@ 2.1 Problem Formulation
Table 1: The sub-tasks of EFSA
@@ 2.2 The Sub-tasks of EFSA
The EFSA task could be further divided into two event-level sub-tasks, namely Coarse-grained EFSA (C-EFSA) and Fine-grained EFSA (F-EFSA), along with an entity-level FSA task. These tasks are outlined in Table 1. Due to the granularity difference of events categorization within different tasks, the difficulty of the three tasks, C-EFSA, F-EFSA, and complete EFSA, exhibits an increasing trend. This enables FSA researchers to conduct experiments on tasks of varying difficulty levels and financial practitioners to analyze market conditions at different event granularities. Additionally, our dataset can also support the existing FSA task. By disregarding event labels, our task can be simplified to a purely entity-level FSA task.
@@ 2.3 Relatedness to ABSA Task
Recall that ABSA aims to identify sentiment elements related to a specific text, which could either be single or multiple elements, including the dependency relations among them (Zhang et al., 2022). These sentiment elements comprise aspect terms, sometimes aspect categories, opinion terms, and aspect-level sentiment. Among them, aspect and opinion terms are specific text spans within a sentence, and identifying them is an extraction task. Determining aspect categories and sentiments are usually classification tasks.
Although EFSA and ABSA appear to be similar in form, as they both involve extracting the subject of sentiment from the original text, i.e. companies and aspect terms, and identifying sentiment-related phrases (events or opinion terms), and finally determining sentiment, EFSA presents two main differences. Firstly, the events in EFSA may be extremely long and discontinuous, which results in our formulation simplifying the event extraction to a classification task. Few existing ABSA approaches can address such change without any modification. Secondly, the financial context makes its mapping from events to sentiments unique in EFSA. The mapping is constructed on domain expertise rather than subjective emotional expressions. These two reasons essentially differentiate EFSA from ABSA.
While the sub-tasks of EFSA exhibit a certain degree of similarity to the ABSA task. For instance, the entity-level FSA can be addressed by most current ABSA approaches. Hence, in subsequent experiments in Section 4.1, we benchmark several applicable ABSA baselines on our dataset.
@@ 3 Dataset
Figure 2: The statistics of data distribution of our dataset. (a) presents a nested pie chart that illustrates the distribution of event labels. The inner layer and the outer layer represent the distribution of coarse- and fine-grained events; the dark and light shades of the same color represent the coarse-grained event and its subdivided fine-grained events. (b), (c), (d), and (e) present the distribution among different publish times, overall sentiments, news body length, and sentiment distribution of each coarse-grained event. (f) presents the distribution of industries, where various colors denote $32$ distinct industries. The size of the color blocks in (a) and (f) represents the data volume.
Financial sentiment indicators are categorized into market-derived sentiment and human-annotated sentiments (Luo et al., 2018). The market-derived sentiment is estimated based on market dynamics such as stock price changes and trading volume, potentially incorporating noise from other sources (Fei et al., 2023). Therefore, we employed manual labeling conducted by experts in the financial field to ensure the creation of a high-quality dataset. Specifically, the data construction process is elaborated in detail from the following three aspects: data collection, annotation, and distribution.
@@ 3.2 Data Annotation
Labeling System. To address the complex spectrum of event types in financial news, we have developed a detailed event taxonomy by professional financial practitioners. This taxonomy comprises seven coarse-grained categories: Financial Affairs, Shareholder Affairs, Stock Affairs, Compliance and Credit, Management Affairs, Business Operations and Financing and Investment, and further extends to $43$ fine-grained categories. The complete event taxonomy is presented in Appendix A. We encourage FSA researchers to annotate or augment existing datasets with additional annotations, according to our publicized labeling system. For sentiment labels, we adopted a three-category sentiment polarity classification consisting of positive, neutral, and negative. Each company was classified into a industry referencing the Shenwan Industry Classification Standard***The Shenwan Industry Classification is a widely used industry classification standard consisting of $32$ industry categories. It was proposed by Shenwan Hongyuan Securities for financial investment management and research. https://www.swsresearch.com/institute_sw/allIndex/downloadCenter/industryType.
Annotation Platform. To streamline the data annotation process, we developed a specialized annotation platform tailored to our task requirements. This platform presents news bodies to annotators as input, enabling them to choose specific text spans within the news article for company labels and to directly assign labels for events, sentiments, and industries within the system.
@@
Labeling Evaluation Metrics. To ensure consistency across various annotations, we tasked multiple annotators with independently labeling the same article set. Following previous work (Hripcsak and Rothschild, 2005; Barnes et al., 2018), we use different metrics to measure the inter-annotator agreement of different annotation tasks. We employ the AvgAgr metric (Wiebe et al., 2005) to evaluate span extraction annotation consistency. The AvgAgr score of Company is $0.65$. For classification annotation consistency score, we employ Fleiss’ Kappa (Fleiss, 1971) to evaluate. The Fleiss’ Kappa scores of classification annotation are as follows: event ($0.62$), sentiment ($0.67$), and industry ($0.70$). The results, which fall between $0.61$ and $0.8$, demonstrate a substantial agreement(Landis JRKoch, 1977) among different annotators.
@@ 3.3 Data Distribution
We balanced the distribution of different elements in our dataset to ensure a proportional data distribution. Specifically, we balanced the following label distributions:
Figure 2 (a) demonstrates the substantial uniformity of coarse- and fine-grained event distribution. For each fine-grained event, there is a sufficient amount of data to support, ensuring no instances no data shortage exists for any particular category. Most of our data’s publication times span from 2021 to 2023. As shown in Figure 2 (b), We balanced the distribution of years to reflect market dynamics evenly across different periods. We restricted the length of the news context to 300 Chinese characters. Besides, we balanced the distribution of context length to manage the complexity of the task (c.f. Figure 2 (d)), with the average length of the news body in our dataset being $148.5$. Furthermore, we balanced the overall sentiment distribution and the sentiment distribution of each coarse-grained event. Figure 2 (c) and (e) show that our dataset achieves a balance in overall sentiment distribution while also ensuring that there are no significant imbalances in sentiment distribution within specific coarse-grained events. For fine-grained events, given that certain events have specific sentiment tendencies, such as Legal Affairs often corresponding to negative emotions, we did not balance the sentiment distribution of each fine-grained event. We also calculated the distribution of industries  (c.f. Figure 2 (f)) to ensure our dataset comprehensively and uniformly covers a wide range of sectors.
@@ 4.1 Benchmark Methods
LLMs. We selected Chinese LLMs in general domain with strong performance and financial domain-specific LLMs. We interacted with the LLMs by constructing prompts to instruct LLMs to generate structured outputs. Our specific prompts are reported in Appendix B. Particularly, we fine-tuned the open-source, deployable general domain LLMs using LoRA (Hu et al., 2021) on our dataset.
@@
(2) ChatGLM (Du et al., 2021) is a bilingual language model, supporting both Chinese and English, built on the General Language Model (GLM) framework. We evaluated Chatglm-3 on our dataset, which achieved the best performance on the Chinese LLM evaluation benchmark C-EVAL (Huang et al., 2023). We further evaluated the performance of the latest ChatGLM-4 released on January 2024 by querying interface glm-4 provided by ZHIPU AI open platform. This method stands out as one of the LLMs known for its robust Chinese language alignment capabilities.
@@
(5) QwenLM (Bai et al., 2023) is a bilingual language model proficient in both Chinese and English, which achieves better performance than LLaMA2-70B on all tasks and outperforms GPT-3.5 on 7 out of 10 benchmarks (Bai et al., 2023).
@@
(7) Xuanyuan (Zhang and Yang, 2023) is a financial domain-specific LLM derived through incremental pre-training based on LLma2. It exhibits significant enhancements in both Chinese language proficiency and financial capabilities.
@@
Leveraging the GAS’s framework (Zhang et al., 2021b), we benchmarked the performance of the mT5-large and BART-large-chinese (Zhang et al., 2021a) models on Entity-level sub-task. We also benchmark E2E-BERT (Li et al., 2019) on this sub-task. We randomly split $80\%$ data for training and the left $20\%$ part is for testing. It is noted that the original training hyperparameters are used for each respective model.
@@ 4.2 Our Framework
Given Chain of Thought (CoT)’s success in analyzing implicit sentiment (Fei et al., 2023) and the inherent chained progression relationship between coarse- and fine-grained events in the EFSA task, we devised a four-hop CoT framework for our task. It involves four steps. Step 1. Utilize a news article as input and instruct LLM to identify the company mentioned within the text. Step 2. Subsequently, use the news article and the identified companies as input. Instruct the LLM to choose a corresponding event from a predefined set of coarse-grained events. Step 3. Use news article and (company, coarse-grained-event) tuple as input. Instruct LLM to choose a corresponding event from a predefined set of fine-grained events. Step 4. Use news article and (company, coarse-grained event, fine-grained event) triplet as input. Instruct LLM to choose a corresponding sentiment of the triplet.
Figure 3 illustrates our four-hop CoT framework. The details of the four-hop CoT prompt are displayed in Appendix B. Based on this framework, we employed dialogue fine-tuning on deployable LLMs to evaluate effectiveness.

### v2

@@ Abstract
In this paper, we extend financial sentiment analysis (FSA) to event-level since events usually serve as the subject of the sentiment in financial text. Though extracting events from the financial text may be conducive to accurate sentiment predictions, it has specialized challenges due to the lengthy and discontinuity of events in a financial text. To this end, we reconceptualize the event extraction as a classification task by designing a categorization comprising coarse-grained and fine-grained event categories. Under this setting, we formulate the Event-Level Financial Sentiment Analysis (EFSA for short) task that outputs quintuples consisting of (company, industry, coarse-grained event, fine-grained event, sentiment) from financial text. A large-scale Chinese dataset containing $12,160$ news articles and $13,725$ quintuples is publicized as a brand new testbed for our task. A four-hop Chain-of-Thought LLM-based approach is devised for this task. Systematically investigations are conducted on our dataset, and the empirical results demonstrate the benchmarking scores of existing methods and our proposed method can reach the current state-of-the-art. Our dataset and framework implementation are available at https://github.com/cty1934/EFSA.
@@
Tianyu Chen1,2,3111These authors contributed equally to this work., Yiming Zhang4111These authors contributed equally to this work., Guoxin Yu1,2,3, Dapeng Zhang5, Li Zeng6222Correspondence to Xiang Ao and Li Zeng., Qing He1,2,3,4, Xiang Ao1,2,3222Correspondence to Xiang Ao and Li Zeng. 1Key Laboratory of AI Safety, Chinese Academy of Sciences (CAS), Beijing, China. 2Key Lab of Intelligent Information Processing, Institute of Computing Technology, CAS, Beijing, China. 3University of Chinese Academy of Sciences, Beijing, China. 4Henan Institute of Advanced Technology, Zhengzhou University, Zhengzhou, China. 5School of IoT Engineering, Jiangsu Vocational College of Information Technology, Wuxi, China. 6Information Technology Department I, Shenzhen Stock Exchange. {chentianyu22s,aoxiang}@ict.ac.cn
@@ 1 Introduction
To this end, we aim to discover the events in financial text to provide an easier financial sentiment prediction. An intuitive way to identify an event from financial text is extracting a specific text span from the original text just like existing aspect-based sentiment analysis (ABSA) tasks that find both aspect and opinion terms from customer’s comments (Zhang et al., 2022b; Yu et al., 2023, 2021b). However, direct adapting approaches of ABSA for this task could be ineffective even infeasible due to the events in the financial text being overlong and discontinuous (c.f. Figure 1). We will provide a further discussion about the similarities and differences between ABSA and our task in Section 2.3.
In this paper, based on our observations, we provide an alternative setting for FSA and propose a novel task, named Event-Level Financial Sentiment Analysis (EFSA), involving the prediction of quintuples (company, industry, coarse-grained event, fine-grained event, sentiment). Here we have enhanced the FSA tasks in two facets. First, to overcome the difficulties associated with extracting events from financial texts, we reconceptualize the event extraction task as a classification task. We design a categorization system that comprises both coarse-grained and fine-grained event categories, specifically tailored for various event types in financial news. Besides, we construct a knowledge-based rule to classify companies by industry, enabling FSA to elevate to higher dimensions, such as indices or sector markets. Our task setting can offer substantial practical value in financial applications, including stock trading, stock market anomaly attribution analysis, enterprise risk management, etc. Figure 1 presents example quintuples in our EFSA task.
To support this task, we annotated a large-scale dataset from Chinese financial news. The dataset includes $12,160$ news articles, selected from an initial set of over $50,000$ articles collected from mainstream Chinese financial news websites. Detailed annotations were conducted based on the above task settings. To the best of our knowledge, this dataset is a large-scale fine-grained annotated dataset for FSA and the largest Chinese dataset in the event-level FSA domain.
@@
  • We conduct systematically benchmark experiments to investigate the efficacy of existing methods for the EFSA task and introduce a novel LLM-based framework reaching the current state-of-the-art.
@@ 2.1 Problem Formulation
Table 1: The sub-tasks of EFSA.
@@ 2.2 The Sub-tasks of EFSA
The EFSA task could be further divided into two event-level sub-tasks, namely Coarse-grained EFSA (C-EFSA) and Fine-grained EFSA (F-EFSA), along with an entity-level FSA task. These tasks are outlined in Table 1. Due to the granularity difference of events categorization within different tasks, the difficulty of the three tasks, C-EFSA, F-EFSA, and complete EFSA, exhibits an increasing trend. This enables FSA researchers to conduct experiments on tasks of varying difficulty levels and financial practitioners to analyze market conditions at different event granularities. Additionally, our dataset can also support existing FSA tasks. By omitting event classification, our task can be simplified to a purely entity-level FSA task.
@@ 2.3 Relatedness to ABSA Task
Recall that ABSA aims to identify sentiment elements related to a specific text, which could either be single or multiple elements, including the dependency relations among them (Zhang et al., 2022b; Yu et al., 2021a). These sentiment elements comprise aspect terms, sometimes aspect categories, opinion terms, and aspect-level sentiment. Among them, aspect and opinion terms are specific text spans within a sentence, and identifying them is an extraction task. Determining aspect categories and sentiments are usually classification tasks.
Although EFSA and ABSA appear to be similar in form, as they both involve extracting the sentiment elements from the original text, i.e. companies and aspect terms, and identifying sentiment-related phrases (events or opinion terms), and finally determining sentiment, EFSA presents two main differences. Firstly, the events in EFSA may be extremely long and discontinuous, which results in our formulation simplifying the event extraction to a classification task. Few existing ABSA approaches can address such change without any modification. Secondly, the financial context makes its mapping from events to sentiments unique in EFSA. The mapping is constructed on domain expertise rather than subjective emotional expressions. These two reasons essentially differentiate EFSA from ABSA.
The sub-tasks of EFSA exhibit a certain degree of similarity to the ABSA task. For instance, the entity-level FSA can be addressed by most current ABSA approaches. Hence, in subsequent experiments in Section 4.1, we benchmark several applicable ABSA baselines on our dataset.
@@ 3 Dataset
Figure 2: The statistics of data distribution of our dataset. (a) presents a nested pie chart that illustrates the distribution of event labels. The inner layer and the outer layer respectively represent the distribution of coarse- and fine-grained events; the dark and light shades of the same color represent the coarse-grained event and its subdivided fine-grained events. (b), (c), (d), and (e) present the distribution among different publish times, overall sentiments, news body length, and sentiment distribution of each coarse-grained event. (f) presents the distribution of industries, where various colors denote $32$ distinct industries. The size of the color blocks in (a) and (f) represents the data size.
Financial sentiment indicators are categorized into market-derived sentiment and human-annotated sentiments (Luo et al., 2018). The market-derived sentiment is estimated based on market dynamics such as stock price changes and trading volume, potentially incorporating noise from other sources (Fei et al., 2023). Therefore, we employed manual labeling conducted by financial experts to ensure the creation of a high-quality dataset. Specifically, the data construction process is elaborated in detail from the following three sections: data collection, data annotation, and data distribution.
@@ 3.2 Data Annotation
Labeling System. To address the complex spectrum of event types in financial news, we developed a detailed event taxonomy by professional financial practitioners. This taxonomy comprises seven coarse-grained categories: Financial Affairs, Shareholder Affairs, Stock Affairs, Compliance and Credit, Management Affairs, Business Operations and Financing and Investment, and further extends to $43$ fine-grained categories. The complete event taxonomy is presented in Appendix A, serving as a labeling reference for researchers. For sentiment labels, we adopted a three-category sentiment polarity classification consisting of positive, neutral, and negative. Each company was classified into a industry referencing the Shenwan Industry Classification Standard***The Shenwan Industry Classification is a widely used industry classification standard consisting of $32$ industry categories. It was proposed by Shenwan Hongyuan Securities for financial investment management and research. https://www.swsresearch.com/institute_sw/allIndex/downloadCenter/industryType.
Annotation Platform. To streamline the data annotation process, we developed a specialized annotation platform tailored to our task requirements. This platform presents news bodies to annotators as input, enabling them to choose specific text spans within the news article for company labels and to directly assign event labels, sentiment labels, and industry labels within the system. The screenshot of the annotation platform is shown in Appendix B.
@@
Labeling Evaluation Metrics. To ensure consistency across various annotations, we tasked multiple annotators with independently labeling the same news article set. Following previous work (Hripcsak and Rothschild, 2005; Barnes et al., 2018), we use different metrics to measure the inter-annotator agreement of different annotation tasks. We employ the AvgAgr metric (Wiebe et al., 2005) to evaluate span extraction annotation consistency. The AvgAgr score of Company is $0.65$. For classification annotation consistency score, we employ Fleiss’ Kappa (Fleiss, 1971) to evaluate. The Fleiss’ Kappa scores of classification annotation are as follows: fine-grained event ($0.62$), sentiment ($0.67$), and industry ($0.70$). The results, which fall between $0.61$ and $0.8$, demonstrate a substantial agreement (Landis JRKoch, 1977) among different annotators.
@@ 3.3 Data Distribution
We balanced the distribution of different elements in our dataset to ensure a proportional data distribution. The specific data distribution is reported as follows:
Figure 2 (a) demonstrates the substantial uniformity of coarse- and fine-grained event distribution. For each fine-grained event, there is a sufficient amount of data to support, ensuring no instances no data shortage exists for any particular category. As shown in Figure 2 (b), We balanced the distribution of years to reflect market dynamics evenly across different periods. Most of our data’s publication times span from 2021 to 2023. Figure 2 (d) demonstrates the distribution of context length. We restricted the length of the news context to 300 Chinese characters. Besides, we balanced the distribution of context length to manage the complexity of the task, with the average length of the news body in our dataset being $148.5$. Figure 2 (c) and (e) demonstrate that our dataset achieves a balanced overall sentiment distribution. Additionally, it maintains a balance between positive and negative sentiments within specific coarse-grained events, since financial texts inherently tend to exhibit a sentiment bias rather than being neutral (Cortis et al., 2017; de França Costa and da Silva, 2018). For fine-grained events, given that certain events have specific sentiment tendencies, such as Legal Affairs often corresponding to negative emotions, we did not balance the sentiment distribution of each fine-grained event. We also calculated the distribution of industries  (c.f. Figure 2 (f)) to ensure our dataset comprehensively and uniformly covers a wide range of sectors.
@@ 4.1 Benchmark Methods
LLMs. We prioritized models with strong Chinese language capabilities, selecting general domain LLMs and financial domain-specific LLMs. Particularly, we fine-tuned the open-source, deployable general domain LLMs using LoRA (Hu et al., 2021) on our dataset. We interacted with the LLMs by constructing prompts to instruct LLMs to generate structured outputs. To ensure fairness in our benchmark tests, we use the same prompts for every LLMs under different settings, which are detailed in Appendix C.
@@
(2) ChatGLM (Du et al., 2021) is a Chinese and English bilingual language model, constructed utilizing the General Language Model (GLM) framework. ChatGLM-3 demonstrated superior performance on the Chinese LLM evaluation benchmark C-EVAL (Huang et al., 2023). Moreover, we extended our analysis to the latest iteration, ChatGLM-4, by querying the API interface glm-4 provided by ZHIPU AI open platform. ChatGLM-4 stands out as one of the LLMs known for its robust Chinese language alignment capabilities.
@@
(5) QwenLM (Bai et al., 2023) is a Chinese and English bilingual language model. It achieved better performance than LLaMA2-70B on all tasks and outperforms GPT-3.5 on 7 out of 10 benchmarks (Bai et al., 2023).
@@
(7) Xuanyuan (Zhang and Yang, 2023) is a financial domain-specific LLM derived through incremental pre-training based on Llama2. It exhibits significant enhancements in both Chinese language proficiency and financial capabilities.
@@
Leveraging the GAS (Generative ABSA)’s framework (Zhang et al., 2021b), we benchmarked the performance of the mT5-large and BART-large-chinese (Zhang et al., 2021a) models on entity-level sub-task. We also benchmarked E2E-BERT (Li et al., 2019) on this sub-task. We randomly split $80\%$ data for training and the left $20\%$ part is for testing. It is noted that the original training hyperparameters are used for each respective model.
@@ 4.2 Our Framework
Given Chain of Thought (CoT)’s success in analyzing implicit sentiment (Fei et al., 2023) and the inherent chained progression relationship between coarse- and fine-grained events in the EFSA task, we devised a four-hop CoT framework for our task. It involves four steps. Step 1. Utilize a news article as input and instruct the LLM to identify the company mentioned within the text. Step 2. Subsequently, use the news article and the identified companies as input. Instruct the LLM to choose a corresponding event from a predefined set of coarse-grained events. Step 3. Use the news article and the (company, coarse-grained-event) tuple as input. Instruct the LLM to choose a corresponding event from a predefined set of fine-grained events. Step 4. Use the news article and the (company, coarse-grained event, fine-grained event) triplet as input. Instruct the LLM to choose a corresponding sentiment of the triplet.
Figure 3 illustrates our four-hop CoT framework. The details of the four-hop CoT prompt are displayed in Appendix C. Based on this framework, we employed dialogue fine-tuning on deployable LLMs to evaluate effectiveness.